OpenStudy (anonymous):

Simplify the given expression: (−5 + 3i) • (1 − 2i)

OpenStudy (anonymous):

This is like the other one. Use foil first.

OpenStudy (anonymous):

okay -5 * 1 -5*-2i 3i*1 3i*-2i

OpenStudy (anonymous):

FOIL means first, inner, outer, last.

OpenStudy (anonymous):

thats what i did .

OpenStudy (anonymous):

-5 * 1 3i * 1 -5 * -2i 3i * -2i

OpenStudy (anonymous):

okayy now what is the next step

OpenStudy (anonymous):

Tell me what you have after multiplying all of those.

OpenStudy (anonymous):

-5 * 1=-5 3i * 1=3i -5 * -2i =10i 3i * -2i= -6i

OpenStudy (anonymous):

When you multiply 3i and -2i you get -6i^2.

OpenStudy (anonymous):

So you have -5 + 3i + 10i - 6i^2 Remember what i^2 is.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

So -5 + 3i + 10i - 6(-1)

OpenStudy (anonymous):

now ?

OpenStudy (anonymous):

Now... solve it.

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

-30 + 10i Im guessing

OpenStudy (anonymous):

ughhh im pretty sure thats wrong because these are my choices −5 − 6i 1 + 13i 2 + i 4 − 5i

OpenStudy (anonymous):

+ 13 I Meant

OpenStudy (anonymous):

-5 + 3i + 10i - 6(-1) Okay, well you're doing it wrong. What is: -6 * -1

OpenStudy (anonymous):

= 6

OpenStudy (anonymous):

Right, so -5 + 3i + 10i - 6(-1) -5 + 3i + 10i + 6

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

13i +1

OpenStudy (anonymous):

1 + 13i. It's better like that. :D

OpenStudy (anonymous):

okayy now i got it im just not sure how 1^2 Comes Into Play

OpenStudy (anonymous):

You're not sure how i^2 = -1?

OpenStudy (anonymous):

no why 1^2 is put in to the equation

OpenStudy (anonymous):

You mean you don't know why i * i = i^2?

OpenStudy (anonymous):

like is it always going to be 1^2

OpenStudy (anonymous):

I don't understand, there is no 1^2 in there.

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

you asked in the beginning what is 1^2 Right I Said -1 So The You Put It In The Equation -5 + 3i + 10i - 6(-1) <------ why

OpenStudy (anonymous):

1^2 is 1. I said i as in the letter i.

OpenStudy (anonymous):

i^2 = -1 Don't question why, because it's confusing and it won't ever help you in real life. Just accept it. :D

OpenStudy (anonymous):

So 6i^2 is 6 * i^2 and i ^2 = -1 so you sub it in and then you have 6 * -1 6(-1)

OpenStudy (anonymous):

Do you get it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, try simplifying this so I can make sure you understand it well. (5 - 2i)(-5 + 4i)